OpenStudy (anonymous):

k/6+2=0 solve each equation.

OpenStudy (anonymous):

k/6 = -2 k = -2*6 k= -12

OpenStudy (anonymous):

thanks harkirat

OpenStudy (anonymous):

solving means that u hv to find the value of the variable.. for this u shift to other sides numbers that r being added to the term with the variable then u shift the number which is dividing the varibale...

OpenStudy (anonymous):

thats my dog!!